In applying the ε-δ definition of continuity to the function f(x)=8x-5 at x=1, given ε>0, we can choose

a) δ=8/ε

b) δ=ε/8

c) δ=5/ε

d) δ=ε/5

e) δ=8ε


1
Expert's answer
2021-02-03T03:58:55-0500

Let "\varepsilon>0" is given. Choose "\delta=\varepsilon/8."

Assume "|x-1|<\delta."

Thus


"|(8x-5)-3|=|8x-8|"

"=|8||x-1"

"=8|x-1|"

"<8\delta"

"=8\cdot\dfrac{\varepsilon}{8}"

"=\varepsilon"


Therefore


"\lim\limits_{x\to1}(8x-5)=3=8(1)-5=f(1)"

b) δ=ε/8